wilma is twice as old as jamal, and jamal is 3 years older than tim. the sum of all their ages is 17. how old is jamal?
Answer by stanbon(75887)

wilma is twice as old as jamal, and jamal is 3 years older than tim. the sum of all their ages is 17. how old is jamal?
-----
Equations:
w = 2j
j = t+3
w + j + t = 17
-----
Substitute and solve for "j":
2j + j + j-3 = 17
4j = 20
j = 5 (Jamal's age)
------------
Cheers,
Stan H.
===============